# Fully heavy tetraquark states with diquark-antidiquark configuration

### v1

@@ III.2.3 $cb][\bar{c}\bar{b}]$ system
There are eight $P$-wave tetraquark states with $J^{PC}=1^{--}$. Their masses and decay channels are exhibited in Table 3. It is worth noting that the state with a mass of 13067.9 MeV lies very close to the $B_{c}^{*}\bar{B}_{c1}$ threshold. As a result, it can easily decay into this channel and is expected to appear as a narrow resonance due to the proximity to the threshold.
@@
There are four $1^{-+}$ states with masses of 12949.7 MeV, 12993.1 MeV, 13019.9 MeV, and 13057.7 MeV. All of these states lie above the $\eta_{c}\chi_{b1}$ and $\chi_{c1}\eta_{b}$ thresholds, and can therefore naturally decay into these two channels. In addition, the state with a mass of 12993.1 MeV also lies above the $h_{c}\Upsilon$ threshold, making this decay channel accessible as well. The two higher-mass states, 13019.9 MeV and 13057.7 MeV, lie above even higher thresholds, including $\chi_{c1}\eta_{b}$ and $B_{c}\bar{B}_{c1}$. Furthermore, the heaviest state, with a mass of 13057.7 MeV, lies above the $B_{c}^{*}\bar{B}_{c0}$ threshold and can decay into this channel as well.
III.3 $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$ and $cc][\bar{b}\bar{b}]$ systems
@@
For the $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$, and $cc][\bar{b}\bar{b}]$ systems, $C$-parity is not well-defined due to their flavor asymmetry. As a result, the possible quantum numbers for these configurations are $J^{P}=0^{+}$, $1^{+}$, $2^{+}$, $0^{-}$, and $1^{-}$. Our theoretical results are shown in Fig. 2 and Tables 4-6. The analysis method is similar to that for the $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$and $cb][\bar{c}\bar{b}]$ tetraquark states.
@@ IV Summary
In this paper, we systematically study the mass spectra of the lower $S$- and $P$-wave fully heavy tetraquark states, as well as analysis their possible strong decay channels. Considering the basic features of the fully heavy tetraquark system, we adopted an improved Hamiltonian based on the one gluon exchange process. The Hamiltonian include constituent quark masses, spin-spin interaction, spin-orbit coupling and pure orbital contribution. The model parameters required for our calculations are derived from fitting conventional mesons and baryons. The parameters corresponding to different color structures can be calculated through $SU(3)$ group algebra.
For the fully charmed tetraquark system, our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as $S$-wave fully charmed tetraquark states. It is worth noting that we predict two fully charmed tetraquark states with quantum numbers $J^{PC}=1^{--}$, whose masses are 6524.51 MeV and 6555.97 MeV respectively. At the same time, we analyzed the decay states of these two states and found that their characteristics are consistent with the observations of the CMS Collaboration on the $X(6600)$ resonance. For the fully bottom tetraquark system, our calculated masses are all above the $\Upsilon\Upsilon$ threshold and exceed the $bb\bar{b}\bar{b}$ mass reported by CMS Collaboration and ANDY Collaboration. Nevertheless, LHCb Collaboration and other experiments have not observed any $bb\bar{b}\bar{b}$ resonant states thus far. Further experimental investigations are required to address this question. In addition, we calculated masses of several other fully heavy tetraquark states and analysis their strong decay channels. It is also anticipated that some of the states predicted in this work may be observed in future experiments.
@@ Acknowledgements.
The authors would like to thank Yanmei Xiao for helpful comments and discussions. This work was supported by the Scientific Research Foundation of Chengdu University of Technology under grant No.10912-KYQD2022-09557.
@@ Appendix A Wave Functions
The wave functions of all fully heavy tetraquark states are given below.

### v2

@@ III.2.3 $cb][\bar{c}\bar{b}]$ system
There are eight $P$-wave tetraquark states with $J^{PC}=1^{--}$. Their masses and decay channels are exhibited in Table 3. Similarly, we can find some states with relatively narrow decay widths, which warrants closer attention in experiments. Of particular note is the fact that the tetraquark state with a mass of 13067.9 MeV is very close to the $B_{c}^{*}\bar{B}_{c1}$ threshold. As a result, it can easily decay into the mesons $B_{c}^{*}\bar{B}_{c1}$. And this state is expected to appear as a narrow resonance due to the proximity to the threshold.
@@
For quantum numbers $1^{-+}$, there are four $cb][\bar{c}\bar{b}]$ states with masses of 12949.7 MeV, 12993.1 MeV, 13019.9 MeV, and 13057.7 MeV. All of these states lie above the $\eta_{c}\chi_{b1}$ and $\chi_{c1}\eta_{b}$ thresholds, and can therefore naturally decay into these two channels. In addition, the state with a mass of 12993.1 MeV also lies above the $h_{c}\Upsilon$ threshold, making this decay channel accessible as well. In particular, the narrow decay width suggests that this state may form a peak in the $h_{c}\Upsilon$ invariant mass spectrum. The two higher-mass states, 13019.9 MeV and 13057.7 MeV, lie above even higher thresholds, including $\chi_{c1}\eta_{b}$ and $B_{c}\bar{B}_{c1}$. Furthermore, the heaviest state, with a mass of 13057.7 MeV, lies above the $B_{c}^{*}\bar{B}_{c0}$ threshold and can decay into this channel as well. Of course, based on decay width analysis, 13019.9 MeV and 13057.7 MeV are more likely to be found in the invariant mass spectra of $B_{c}\bar{B}_{c1}$ and $B_{c}^{*}\bar{B}_{c0}$, respectively.
III.3 $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$, and $cc][\bar{b}\bar{b}]$ systems
@@
For other fully heavy tetraquark $cc][\bar{c}\bar{b}]$, $bb][\bar{b}\bar{c}]$, and $cc][\bar{b}\bar{b}]$ systems, $C$-parity cannot be well defined due to the violation of flavor symmetry in quark components. As a result, the possible quantum numbers for these configurations are $J^{P}=0^{+}$, $1^{+}$, $2^{+}$, $0^{-}$, and $1^{-}$. Our theoretical results are shown in Fig. 2 and Tables 4-6. The analytical method is similar to the previous tetraquark $cc][\bar{c}\bar{c}]$, $bb][\bar{b}\bar{b}]$, and $cb][\bar{c}\bar{b}]$ systems. Therefore, we will not elaborate on it here.
@@ IV Summary
In this paper, we systematically study the mass spectra of the lower $S$- and $P$-wave fully heavy tetraquark states, as well as analysis their possible strong decay channels. Considering the basic features of the fully heavy tetraquark system, we adopted an improved Hamiltonian based on the one gluon exchange process. The Hamiltonian exhibits a diquark-antidiquark configuration, specifically including effective mass of constituent quark, spin-spin interaction, spin-orbit coupling, and pure orbital contribution. The main model parameters required for our calculations are derived from fitting conventional mesons and baryons. Other parameters $\kappa_{ij}$ corresponding to different color structures can be calculated through $SU(3)$ group algebra.
For the study of fully charmed tetraquark systems, our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as $1S$-wave fully charmed tetraquark states. It is worth noting that we predict a $1S$-wave fully charmed tetraquark state $cc\bar{c}\bar{c}$ with quantum numbers $J^{PC}=2^{++}$, its mass is 6260.89 MeV. Meanwhile, we analyzed the decay channels of this tetraquark state and further confirmed that its characteristics are consistent with the observational results of the $X(6200)$ resonance found in the recent study of the LHCb dataset. For the fully bottom tetraquark system, our calculated mass spectra are all above the $\Upsilon\Upsilon$ threshold and exceed the $bb\bar{b}\bar{b}$ mass reported by CMS Collaboration and ANDY Collaboration. Nevertheless, LHCb Collaboration and other experiments have not observed any $bb\bar{b}\bar{b}$ resonant states thus far. In addition, we calculated masses of several other fully heavy tetraquark states and analysis their strong decay channels. Finally, it is anticipated that some of the narrower fully heavy tetraquark states predicted in this work can be observed in future experiments.
@@ Acknowledgements.
The authors would like to thank Yanmei Xiao for helpful comments and discussions. This work was supported by the Scientific Research Foundation of Chengdu University of Technology under grant No.10912-KYQD2025-09557.
@@ Appendix A Wave Functions
The basis wave functions of all fully heavy tetraquark states are given below.